OpenStudy (anonymous):

Jocelyn has a jar of 54 coins consisting of dimes and quarters. If she has a total of $9.60, how many dimes does she have?

OpenStudy (anonymous):

\[25x+10y=960\]\[x+y=54\] would be one way to solve

OpenStudy (anonymous):

put \(x\) as the number of quarters, then since the total number of coins is 54 the number of dimes you have is \(54-x\) then \[25x+10(54-x)=960\] solve for \(x\)

OpenStudy (anonymous):

you got it from there?

OpenStudy (anonymous):

ummok let me solve it

OpenStudy (anonymous):

I got 284.8

OpenStudy (anonymous):

is it 28 coins?